2.1. How is the hatching of water python eggs influenced by the temperature of the snakes nest? Researchers placed 104 newly laid eggs in a hot environment, 56 in a neutral environment, and 27 in a cold environment. Hot duplicates the warmth provided by the mother python. Neutral and cooler are cooler, as when the mother is absent. The results: 75 of the hot eggs hatched, along with 38 of the neutral eggs and 16 of the cold eggs. Make a two-way table of environment temperature against hatched or not. .The researchers anticipated that eggs would hatch less well at cooler temperatures. Do the data support that anticipation?
Answer #1

a)

two way table

hot neutral cold total
hatched 75 38 16 129
no hatch 29 18 11 58
total 104 56 27 187

b)

P(hatching| cold) = 16/27 = 59.26%

P(hatching|neutral)=38/56=67.86%

P(hatching|hot)=75/104=72.11%

since, Probability of hatching in cold is least,

so, yes, data support the antipication